OpenStudy (unklerhaukus):

\[\begin{array} &\int 1&\pm2 \\1^{x^z} & \varrho &\langle x\rangle&i&\sum\Sigma \\4\mathcal {F4} &\mathcal{X}&\chi& \mathcal{L}&\\ &\alpha& \Gamma&\\&500&\infty \%\\100_{\lambda} \\\cdots &\Psi^\dagger\Psi\dots \ddots &\vdots&\end{array}\]

sam (.sam.):

nice

OpenStudy (unklerhaukus):

\[\begin{array}\\ \times& \times &\cancel \circ\\ & \cancel \circ &\circ\\ \cancel \circ&\times&\times\end{array}\]